OpenStudy (lollylau):

[Warm-up for previous challenge] Mark only has $1 coins and $2 coins in his wallet (poor guy). He has 10 coins and the total calue of the coins is $16. Find the number of $1 and $2 coins.

OpenStudy (anonymous):

Let the number of $1 coins be x. Number of $2 coins = 10 - x. x + 2(10 - x) = 16 x + 20 - 2x = 16 -x = -4 x = 4 10 - x = 6

OpenStudy (cornzyblack):

hope u understand this the total sum of coins in mark's wallet is 10. But in mark's wallet he has an unknown number of $1 and 2$ coins. And the sum of the total coins in his wallet is $16. So lets take the unknown number of 1$ as x. And the unknown number of 2$ as y. Since the total number of coins in his wallet is 10 hence the sum of the unknown number of coins is also 10 that is, x+y = 10 -------- equation. 1 So then the total value of coins is for simplicity sake i removed the $ sign x(1) + (2)y = 16. ------ equation 2 if x + y = 10 then, x = 10 - y ---- equation 3 so substitute in equation 2, (10 - y) 1 + 2 (y) = 16 10 -y + 2 = 16 12 - y = 16 y = 16 - 12 y = 4 from eqn 1 x + y = 10 substitute the value for x x + 4 = 10 x = 6 so mark has six 2$ coins and four 1$